OpenStudy (anonymous):

solve 8^(x-2)=3 im learning about logarithms if that helps...

OpenStudy (saifoo.khan):

\[\Large a^b = c \] \[\Large b = \log_ac \]

OpenStudy (anonymous):

so x-2=log(8)3?

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

ok thank you!